OpenStudy (anonymous):

The area of a circle is 113.04 in. ^2. What is the radius of the circle?

Directrix (directrix):

A = pi r^2 113.04 = 3.14 r^2 113.04/3.14 = r^2 36 = r^2 6 = r Note: -6 is also an algebraic solution but is not applicable here.